The voltage for the following cell is +0.731 V. Find Kb for the organic base RNH2.

Use E(SCE)= 0.241V.

Pt(s)│H2(1.00 atm)│RNH2(0.100M), RNH3+(0.0500M)║SCE
Answer #1

For the given reaction -

                 R-NH3+ (aq) <----------> R-NH2 (aq) + H+ (aq)

Since - Ecell = -(0.0592/n) log Ka

Where, Ecell = 0.731 - 0.241 = 0.490 V

Therefore, 0.490 = -(0.0592/1) log Ka

Therefore, Ka = 5.30*10-9

Thus, Kb = Kw/ Ka = (1.0*10-14 / 5.30*10-9) = 1.892*10-6